OpenStudy (anonymous):

Apples sell for $1.90 per pound, and bananas sell for $0.75 per pound. Troy bought some apples and some bananas. Together they weighed 3.8 pounds, and cost $5.84. Let a represent the number of pounds of apples Troy purchases. Which expression represents the number of pounds of bananas he purchases? A. 3.8 – a B. a – 3.8 C. 3.8a D. a + 3.8

OpenStudy (anonymous):

You know the total, and you know how many pounds of apples there were.... so how would you figure out how many pounds of bananas?

OpenStudy (anonymous):

dont know

OpenStudy (anonymous):

If there was 1 pound of apples how many pounds of bananas would there be?

OpenStudy (anonymous):

3.8

OpenStudy (anonymous):

Wrong. Try again.

OpenStudy (anonymous):

2.8

OpenStudy (anonymous):

Right. How did you figure that out?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

3.8 – a

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Carrots sell for $2.10 per pound, and crackers sell for $2.90 per pound. Glen bought some carrots and some crackers. The total weight was 2.3 pounds and cost $6.03. Let p represent the number of pounds of carrots. Which equation represents the situation described? A. –0.8p + 6.03 = 6.67 B. 2.1p + 2.9 = 6.03 C. –0.8p + 6.67 = 6.03 D. 2.1p + 6.03 = 2.9

OpenStudy (anonymous):

Ok, this is similar, but slightly more complicated because you're going to have to combine like terms.

OpenStudy (anonymous):

ok so how do i do this

OpenStudy (anonymous):

Ok, if carrots sell for 2.10 a pound and you bought \(p\) pounds of carrots, what's the total cost you paid for carrots?

OpenStudy (anonymous):

i dont know i am in 5th grade

OpenStudy (anonymous):

This is just multiplication. Any 5th grader can do that.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

i am an idiot

OpenStudy (anonymous):

No you're not! You can do this.